# Topological Phases in Non-Hermitian Nonlinear-Eigenvalue Systems

### v1

@@ III Real-band topological phases in non-Hermitian nonlinear systems
[…]
[width=]Fig3.pdf
Figure 3: (a) Band structure of $\omega$ in different $w$. The real and complex edge-mode eigenvalues are marked by red and blue lines, respectively. (b) Band structures of the real part of complex $\lambda$ (blue dots) and real $\lambda$ (orange dots) when $w=4v$. (c) Winding numbers in different $w$. Phase diagrams for the real (d) and complex (e) edge modes. We use $\mathcal{M}=-0.5v^{-2}$, $\gamma=v^{-1}$, and $L=100$.
@@ IV Complex-band topological phases in non-Hermitian nonlinear systems
We study a third-order nonlinear non-Hermitian SSH model. It possesses the same Hermitian Hamiltonian matrix $H_{0}$ as the preceding sections and a non-Hermitian overlap matrix $\mathcal{S}(\omega)=s_{0}+\tilde{s}_{1}\omega+\tilde{s}_{2}\omega^{2}$, where $s_{0}=1$, $\tilde{s}_{1}=-\gamma\sum_{n=1}^{L}(c_{n,A}^{\dagger}c_{n,B}-c_{n,B}^{\dagger}c_{n,A})$ and $\tilde{s}_{2}=-i\mathcal{M}$ [73, 70]. Here, $\gamma$ is a non-reciprocal parameter between the two directions of the nonlinear intracell hopping rates and $\mathcal{M}$ is the nonlinear loss-gain parameter. Equation (1) is formally rewritten as [78]
@@
It shows that the presence of $\bar{s}_{2}$ turns the system into a six-band model. The momentum-space pencil matrix of the auxiliary system is derived to be $P(k,\omega)=\sum_{j=0,\pm}p_{j}(k,\omega)\sigma_{j}$, where $p_{0}(k,\omega)=i\mathcal{M}\omega^{3}-\omega$ and $p_{\pm}(k,\omega)=v\pm\gamma\omega^{2}+we^{\mp ik}$. $P(k,\omega)$ satisfies the chiral symmetry $\Gamma P(k,\omega)\Gamma^{-1}=-P(k,-\omega)$. Its bands touch under the condition $i\mathcal{M}\omega^{3}-\omega=0$, whose solutions are $\omega=0$ and $\pm(-1)^{3/4}\mathcal{M}^{-1/2}$. Therefore, the system may host both the real edge modes at $\omega=0$ and the complex edge modes at $\omega=\pm(-1)^{3/4}\mathcal{M}^{-1/2}$. For the real edge modes at $\omega=0$, the pencil matrix reduces to $P_{\rm R}(k,0)=(v+we^{-ik})\sigma_{+}+\text{H.c.}$, which is just a linear SSH model. At the high-symmetry points $k=0$ and $\pi$, the real bands of $P_{\rm R}(k,0)$ touch at $|v|=|w|$. The pencil matrix for the complex edge modes at $\omega=\pm(-1)^{3/4}\mathcal{M}^{-1/2}$ becomes $P_{\rm C}(k,\pm(-1)^{3/4}\mathcal{M}^{-1/2})=(v-{i\gamma\over\mathcal{M}}+we^{-ik})\sigma_{+}+\text{H.c.}$, whose bands touch at $|k|=|\arccos[\pm v/\sqrt{(\gamma/\mathcal{M})^{2}+v^{2}}]|$ when $w^{2}=(\gamma/\mathcal{M})^{2}+v^{2}$. It is interesting to find that, although the overlap matrix contains non-Hermitian term and the edges modes possess complex eigenvalues, the pencil matrix of the associated auxiliary system is Hermitian. Therefore, it is naturally expected that the original and the auxiliary systems obey the BBC. The winding numbers for the real and complex bands can be defined in $P_{\rm R}(k,0)$ and $P_{\rm C}(k,\pm(-1)^{3/4}\mathcal{M}^{-1/2})$ in a manner similar to that of the Hermitian case.
@@
By numerically solving Eq. (6), we obtain the complex-$\omega$ bands under the OBC in different $w$ [see Fig. 3(a)]. It is remarkable to find that the gapped edge states have not only a real eigenvalue $\omega=0$ (red lines) but also two complex eigenvalues $\omega=\pm(-1)^{3/4}\mathcal{M}^{-1/2}$ (blue lines). The former ones with $\omega=0$ are present when $|w|>|v|$. The latter ones with $\omega=\pm(-1)^{3/4}\mathcal{M}^{-1/2}$ are present when $w^{2}>(\gamma/\mathcal{M})^{2}+v^{2}$. Figure 3(b) shows the $\lambda$ band of the auxiliary system under the OBC. It clearly indicates that a gapless complex edge mode intersects with $\lambda=0$ just at $\omega=0$ and $\pm(-1)^{3/4}\mathcal{M}^{-1/2}$, corresponding exactly to the two types of edge modes in Fig. 3(a). The orange points in the bulk bands reveal that all the values of $\lambda$ are real, whose positions exactly match $\omega=0$ and $\pm(-1)^{3/4}\mathcal{M}^{-1/2}$. Therefore, the auxiliary system at these three positions is Hermitian and we can calculate the topological numbers defined in the pencil matrices $P_{\rm R}(k,0)$ and $P_{\rm C}(k,\pm(-1)^{3/4}\mathcal{M}^{-1/2})$ to characterize the real and complex edge modes, respectively. The obtained winding numbers correctly count the pair number of the two types of edge states [see Fig. 3(c)]. The phase diagrams for the real and complex edge bands in the $\gamma$-$w$ space are depicted in Figs. 3(e) and 3(d). They give a global picture on the non-Hermitian topological phases of this third-order nonlinear-eigenvalue system. To the best of our knowledge, such a topological phase in the complex bands has never been reported before.
@@ V Discussion and Conclusion
Nonlinear-eigenvalue problems arise in various physical systems. For example, in the study of higher-order nonlinear responses of graphene [61, 81, 82], the wave propagation problem is reduced to an eigenvalue problem. Additionally, when analyzing waveguide filter structures with frequency-dependent material properties or specific boundary conditions, their permittivity or permeability becomes frequency-dependent rather than constant [55, 56, 57, 58, 59, 60, 61, 62]. Consequently, the analysis of their resonant characteristics can be formulated as a nonlinear-eigenvalue problem. Meanwhile, waveguide arrays can also be employed to simulate topological insulators [83, 84], including non-Hermitian topological insulators [85, 86, 87]. In the field of structural acoustics, its dynamic equation takes the form of a quadratic nonlinear-eigenvalue problem [88]. Furthermore, acoustic systems can be used similarly to model topological insulators [89, 90, 91, 92, 93]. Therefore, our system can be conceptually derived from the design principles of these practical systems.
@@
VI Acknowledgments
The work is supported by the National Natural Science Foundation of China (Grants No. 124B2090, No. 12275109, No. 92576202, and No. 12247101), the Quantum Science and Technology-National Science and Technology Major Project (Grant No. 2023ZD0300904), the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2025-jdzx07), and the Natural Science Foundation of Gansu Province (No. 22JR5RA389 and No. 25JRRA799).

### v2

@@ III Real-band topological phases in non-Hermitian nonlinear systems
Figure 3: (a) Band structure of $\omega$ in different $w$. The real and complex edge-mode eigenvalues are marked by red and blue lines, respectively. (b) Distribution of the real- (the red line) and complex-band (blue line) edge states and the bulk state (gray line). (c) Band structures of the real part of complex $\lambda$ (blue dots) and real $\lambda$ (orange dots) when $w=4v$. (d) Winding numbers in different $w$. Phase diagrams for the real (e) and complex (f) edge modes. We use $\mathcal{M}=-0.5v^{-2}$, $\gamma=v^{-1}$, and $L=100$.
@@ IV Complex-band topological phases in non-Hermitian nonlinear systems
We study a third-order nonlinear non-Hermitian SSH model. It possesses the same $H_{0}$ as above but with $U_{0}=0$ and a non-Hermitian $\mathcal{S}(\omega)=s_{0}+\tilde{s}_{1}\omega+\tilde{s}_{2}\omega^{2}$, where $s_{0}=1$, $\tilde{s}_{1}=-\gamma\sum_{n=1}^{L}\left(c_{n,A}^{\dagger}c_{n,B}-c_{n,B}^{\dagger}c_{n,A}\right)$ and $\tilde{s}_{2}=-i\mathcal{M}$ [77, 73]. Here, $\gamma$ is a non-reciprocal parameter between the two directions of the nonlinear intracell hopping rates and $\mathcal{M}$ is the nonlinear loss-gain parameter. Equation (1) is rewritten as [82]
@@
The presence of $\tilde{s}_{2}$ turns the system into a six-band model. The pencil matrix is derived to be $P(k,\omega)=\sum_{j=0,\pm}p_{j}(k,\omega)\sigma_{j}$, where $p_{0}(k,\omega)=i\mathcal{M}\omega^{3}-\omega$ and $p_{\pm}(k,\omega)=v\pm\gamma\omega^{2}+we^{\mp ik}$. $P(k,\omega)$ satisfies the chiral symmetry $\mathcal{C}P(k,\omega)\mathcal{C}^{-1}=-P(k,-\omega)$. Its bands touch under the condition $i\mathcal{M}\omega^{3}-\omega=0$, which gives $\omega=0$ and $\pm(-1)^{3/4}\mathcal{M}^{-1/2}$. Thus, the system hosts both the real edge modes at $\omega=0$ and the complex edge modes at $\omega=\pm(-1)^{3/4}\mathcal{M}^{-1/2}$. For the former, the pencil matrix reduces to $P_{\rm R}(k,0)=\left(v+we^{-ik}\right)\sigma_{+}+\text{H.c.}$, which is just a linear SSH model. At the high-symmetry points $k=0$ and $\pi$, the real bands of $P_{\rm R}(k,0)$ touch at $|v|=|w|$. The pencil matrix for the complex edge modes becomes $P_{\rm C}\left[k,\pm(-1)^{3/4}\mathcal{M}^{-1/2}\right]=\left(v-{i\gamma\over\mathcal{M}}+we^{-ik}\right)\sigma_{+}+\text{H.c.}$, whose bands touch at $|k|=|\arccos[\pm v/\sqrt{(\gamma/\mathcal{M})^{2}+v^{2}}]|$ when $w^{2}=(\gamma/\mathcal{M})^{2}+v^{2}$. It is interesting to find that, although the system is non-Hermitian and the edge modes possess complex eigenvalues, the two pencil matrices $P_{\text{R}}(k,0)$ and $P_{\text{C}}\left[k,\pm(-1)^{3/4}\mathcal{M}^{-1/2}\right]$ of the auxiliary system are Hermitian. Therefore, the original and the auxiliary systems obey the BBC. The winding numbers for the real and complex bands can be defined in $P_{\rm R}(k,0)$ and $P_{\rm C}\left[k,\pm(-1)^{3/4}\mathcal{M}^{-1/2}\right]$ in a manner similar to that of the Hermitian case. Without playing any role in the real-band topological phase, the nonreciprocity $\gamma$ and the gain/loss $\mathcal{M}$ together contribute to the complex-band topological phases. $\mathcal{M}$ gives the eigenvalues $\omega=\pm(-1)^{3/4}\mathcal{M}^{-1/2}$ of the topological edge states, while $\gamma$ impacts the critical point of the topological phase transition and the probability distribution of the edge states. First, the Hermiticity of $P_{\text{C}}\left[k,\pm(-1)^{3/4}\mathcal{M}^{-1/2}\right]$ means that the complex-band edge modes have the same BBC as the ones in the Hermitian systems. Second, well protected by the band gaps and chiral symmetry, these modes are robust to the fluctuation of the system parameters. Third, possessing complex eigenvalues, they are dynamically unstable.
@@
By numerically solving Eq. (6), we obtain the complex-$\omega$ bands under the OBC in different $w$ [see Fig. 3(a)]. It is remarkable to find that the gapped edge states have not only a real eigenvalue $\omega=0$ (red lines) but also two complex eigenvalues $\omega=\pm(-1)^{3/4}\mathcal{M}^{-1/2}$ (blue lines). The former ones are present when $|w|>|v|$. The latter ones are present when $w^{2}>(\gamma/\mathcal{M})^{2}+v^{2}$. The distribution of the bulk state with complex eigenvalue in Fig. 3(b) shows the bulk states also reside in the edge, indicating the non-Hermitian skin effect. Figure 3(c) shows the $\lambda$ band of the auxiliary system under the OBC. It clearly indicates that a gapless complex edge mode intersects with $\lambda=0$ just at $\omega=0$ and $\pm(-1)^{3/4}\mathcal{M}^{-1/2}$, corresponding exactly to the two types of edge modes in Fig. 3(a). The orange points in the bulk bands reveal that all the values of $\lambda$ are real, whose positions exactly match $\omega=0$ and $\pm(-1)^{3/4}\mathcal{M}^{-1/2}$. Thus, the auxiliary system at these three positions is Hermitian and we can calculate the topological numbers defined in $P_{\rm R}(k,0)$ and $P_{\rm C}(k,\pm(-1)^{3/4}\mathcal{M}^{-1/2})$ to characterize the real and complex edge modes, respectively. The obtained winding numbers correctly count the pair number of the two types of edge states [see Fig. 3(d)]. The phase diagrams for the real and complex edge bands in the $\gamma$-$w$ space in Figs. 3(e) and 3(f) give a global picture of the non-Hermitian topological phases of this third-order nonlinear-eigenvalue system.
@@ V Discussion and conclusion
Although progress on topological phases in nonlinear-eigenvalue systems is rare, various systems, especially metamaterials, show great potential for realizing our predictions. Nonlinear-eigenvalue problems arise in various physical systems. In the study of higher-order nonlinear responses of graphene [64, 85, 86], the wave propagation problem is reduced to an eigenvalue problem. In analyzing waveguide filter structures with frequency dependent material properties or specific boundary conditions, the frequency-dependence of their permittivity or permeability make the analysis of their resonant characteristics to be a nonlinear-eigenvalue problem [58, 59, 60, 61, 62, 63, 64, 65]. Furthermore, waveguide arrays can also be used to simulate topological insulators [87, 88], including non-Hermitian topological insulators [89, 90, 91]. In the field of structural acoustics, the dynamic equation takes the form of a quadratic nonlinear-eigenvalue problem [92]. Acoustic systems can be used similarly to model topological insulators [93, 94, 95, 96, 97]. Therefore, our system can be conceptually derived from the design principles of these practical systems.
@@
VI Acknowledgment
The work is supported by the National Natural Science Foundation of China (Grants No. 124B2090, No. 12275109, No. 92576202, and No. 12247101), the Quantum Science and Technology-National Science and Technology Major Project (Grant No. 2023ZD0300904), the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2025-jdzx07), and the Natural Science Foundation of Gansu Province (Grants No. 22JR5RA389 and No. 25JRRA799).